# Epistemological Bias As a Means for the Automated Detection of Injustices in Text

### v1

@@ Appendix A Appendix
We scale the inputs to our framework from an initial 20 article headlines used in the comparative test by gathering 1645 articles from the Bing API (up to 100 articles for each topic from no more than 1 month back). We searched the following topics: "Kate Middleton", "Meghan Markle", "Kate Middleton family", "Meghan Markle family", "Kate Middleton outfit", "Meghan Markle outfit", "Kate Middleton child", "Meghan Markle child", "Kate Middleton charity", "Meghan Markle charity", "Kate Middleton wedding", "Meghan Markle wedding", "Kate Middleton moving", "Meghan Markle moving", "Kate Middleton bullied", "Meghan Markle bullied", "Kate Middleton dress", "Meghan Markle dress", "Kate Middleton book", "Meghan Markle book", "Kate Middleton makeup", "Meghan Markle makeup","Kate Middleton health", "Meghan Markle health", "Kate Middleton photo", "Meghan Markle photo", "Kate Middleton money", "Meghan Markle money".
Figure 5: Process of Getting Data to the Framework
In Figure 6 we have a screenshot of the UI shown immediately after a user has submitted a text for analyzing. The word suspected of causing injustice is highlighted with the certainty score, an explanation of why the word was tagged. When the user clicks Show Details they will see more details as displayed in Figure 2.
Figure 6: The interactive interface showing the highlighted tagged word and associated information for a sentence.
In Figure 7, we have a diagram showing how the tagger model computes logits $y_{i}$ using discrete feature $f_{i}$ and BERT embedding $b_{i}$ Pryzant et al. (2020b).
Figure 7: The tagger model computes logits $y_{i}$ using discrete feature $f_{i}$ and BERT embedding $b_{i}$
@@
Table 3: Comparative Headline Sentences Used
Table 4 in this section shows the full output of tagging each headline sentence used in the comparative test. The No. column corresponds to the sentence number, Subject refers to the person subject the headline relates to, Taggerout Bias represents the tagged word from the sentence that could potentially be biased, Taggerout Prob refers to the probability of the tagged word been biased, and Taggerout in lexicon refers to the results of checking whether the corresponding tagged word is observed in the lexicon of associated epistemological biased words. This offers some explanability for the potentially biased word. A True value for this column indicates that the word is indeed in the lexicon, whilst a False value indicates otherwise. Bias Type indicates what kind of epistemological bias type is associated with the tagged word. A regular bias type word indicates that the tagged word was not found in the epistemological bias lexicon. The Sentence Sentiment columns represents the overall sentiment associated with the headline sentence (which can be obtained using any NLP processing tool e.g. spacy).
(a) Positive and Negative Sentimental Context - Scaled Comparative Test. Note the epistemological bias types are abbreviated as such for spacing concerns - Assert: Assertives, Ent: Entailments, Hed: Hedges, Fact: Factives, Neg: Negative, Pos: Positive, Sub: Subjective.
(b) Neutral Sentimental Context - Scaled Comparative Test
Figure 8: Epistemological Bias Types - Scaled Data Comparative Test.
@@
Table 4: Table showing the output of passing each headline through the model that tags the potentially biased word in the given sentence.
@@
No. Stereotype (S) Distance(S) Stereotype Concept(SC) Distance(SC) 1 personal spending habits 0.3457 women should spend money on clothes 0.4914 2 women should be dressed like brides 0.3259 women are property 0.2278 3 feminism 0.3663 sexism 0.2523 4 british women are marginalized for a joke 0.2083 pregnancy 0.3918 5 N/A N/A N/A N/A 6 family law 0.3244 arab folks are not protected 0.2951 7 arab folks should live in a constant state of worry 0.1618 racial hierarchy 0.0999 8 gender hierarchy 0.1178 women are often subjected to terms like "babies" and "pregnant" 0.2632 9 women are sexual objects 0.2602 misogyny 0.2574 10 women should wear dresses 0.3131 women should wear dresses 0.3131 11 women are property of men 0.1226 women are property 0.1494 12 women are sex objects 0.1676 women are vulnerable 0.2495 13 sexual abuse 0.0746 saudi arabians are pedophiles 0.1226 14 sexual assault 0.2091 royalty house guests are paedophiles 0.4691 15 women are sex objects 0.1637 women are property 0.1504 16 black folks want to marry rich people 0.2266 racial marriage 0.0869 17 women are property 0.2026 women are secret business partners 0.3732 18 royalty housekeeping are gold diggers 0.3741 royalty house queens are gold diggers 0.3657 19 sexism 0.1927 gender hierarchy 0.2284 20 women should dress modestly 0.3840 women should wear dresses that cover themselves 0.4047
@@
Table 5: Table showing the potential top ranked stereotype and stereotype concept. We passed each sentence through the CO-STAR and SBF model. After which the potential stereotype and stereotype concepts where semantically ranked to the sentence itself

### v2

@@ Appendix A Appendix
In Figure 9, we have a diagram showing how the tagger model computes logits $y_{i}$ using discrete feature $f_{i}$ and BERT embedding $b_{i}$ (Pryzant et al., 2020b).
Figure 9. The tagger model computes logits $y_{i}$ using discrete feature $f_{i}$ and BERT embedding $b_{i}$
@@
Table 5. Comparative Headline Sentences Used
Table 6 in this section shows the full output of tagging each headline sentence used in the comparative test. The No. column corresponds to the sentence number, Subject refers to the person subject the headline relates to, Taggerout Bias represents the tagged word from the sentence that could potentially be biased, Taggerout Prob refers to the probability of the tagged word been biased, and Taggerout in lexicon refers to the results of checking whether the corresponding tagged word is observed in the lexicon of associated epistemic biased words. This offers some explanability for the potentially biased word. A True value for this column indicates that the word is indeed in the lexicon, whilst a False value indicates otherwise. Bias Type indicates what kind of epistemic bias type is associated with the tagged word. A regular bias type word indicates that the tagged word was not found in the epistemic bias lexicon. The Sentence Sentiment columns represents the overall sentiment associated with the headline sentence (which can be obtained using any NLP processing tool e.g., spacy).
@@
Table 6. Table showing the output of passing each headline through the model that tags the potentially biased word in the given sentence.
@@
No. Stereotype (S) Distance(S) Stereotype Concept(SC) Distance(SC) 1 personal spending habits 0.3457 women should spend money on clothes 0.4914 2 women should be dressed like brides 0.3259 women are property 0.2278 3 feminism 0.3663 sexism 0.2523 4 british women are marginalized for a joke 0.2083 pregnancy 0.3918 5 N/A N/A N/A N/A 6 family law 0.3244 arab folks are not protected 0.2951 7 arab folks should live in a constant state of worry 0.1618 racial hierarchy 0.0999 8 gender hierarchy 0.1178 women are often subjected to terms like “babies” and “pregnant” 0.2632 9 women are sexual objects 0.2602 misogyny 0.2574 10 women should wear dresses 0.3131 women should wear dresses 0.3131 11 women are property of men 0.1226 women are property 0.1494 12 women are sex objects 0.1676 women are vulnerable 0.2495 13 sexual abuse 0.0746 saudi arabians are pedophiles 0.1226 14 sexual assault 0.2091 royalty house guests are paedophiles 0.4691 15 women are sex objects 0.1637 women are property 0.1504 16 black folks want to marry rich people 0.2266 racial marriage 0.0869 17 women are property 0.2026 women are secret business partners 0.3732 18 royalty housekeeping are gold diggers 0.3741 royalty house queens are gold diggers 0.3657 19 sexism 0.1927 gender hierarchy 0.2284 20 women should dress modestly 0.3840 women should wear dresses that cover themselves 0.4047
@@
Table 7. Table showing the potential top ranked stereotype and stereotype concept. We passed each sentence through the CO-STAR and SBF model. After which the potential stereotype and stereotype concepts where semantically ranked to the sentence itself
Headline Participant’s choices (# of participants) Framework Component’s choice (confidence)
Meghan Markle spent a staggering £38,000 on her clothes
for a charity trip
Tagged word: staggering (20)
epibias_type: subjective
sentiment: negative (15)
stereotype: personal spending habits (18)
Injustice Present: Character (14)
Alternative word: significant (3)
Tagged word: staggering (0.999498)
epibias_type: subjective
sentiment: negative
stereotype: personal spending habits
Kate Middleton’s £100,000 Astonishing value of the dress that
won a Prince’s heart (and has hung in a wardrobe for eight years)
Tagged word: astonishing (20)
epibias_type: subjective
sentiment: negative (15)
stereotype: personal spending habits (18)
Injustice Present: Character (14)
Alternative word: significant (3)
Tagged word: astonishing (0.999342)
epibias_type: positively subjective
sentiment: neutral
stereotype: women should be dressed like brides
Meghan Markle’s beloved avocado linked to human rights abuse
and drought, millennial shame
Tagged word: beloved (18)
epibias_type: positively subjective
sentiment: negative (19)
stereotype: feminist are hypocrites (11)
Injustice Present: Character (13)
Alternative word: “ ” (2)
Tagged word: beloved (0.997946)
epibias_type: positively subjective
sentiment: negative
stereotype: feminism
Kate’s morning sickness cure? Prince William gifted with
an avocado for pregnant Duchess
Tagged word: cure (8), gifted (8)
epibias_type: entailments, positively subjective
sentiment: neutral (10)
stereotype: British women are marginalized for a joke (11)
Injustice Present: Character (8)
Alternative word: given (4)
Tagged word: gifted (0.877285)
epibias_type: positively subjective
sentiment: neutral
stereotype: British women are marginalized for a joke
Kate and William ‘packed up the kids’ in search of
‘privacy’ at new Windsor estate
Tagged word: privacy (8)
epibias_type: regular
sentiment: negative (9)
stereotype: children are targets of pedos (17)
Injustice Present: Character (11)
Alternative word: seclusion (2)
Tagged word: packed (0.478948)
epibias_type: regular
sentiment: neutral
stereotype: “ ”
Prince Harry and Duchess Meghan Think Moving to
Canada Will Give Archie a Normal Upbringing
Tagged word: normal (15)
epibias_type: subjective
sentiment: neutral (10)
stereotype: racial hierarchy (15)
Injustice Present: Character (9)
Alternative word: ordinary (2)
Tagged word: normal (0.689049)
epibias_type: subjective
sentiment: neutral
stereotype: racial hierarchy
Not long to go! Pregnant Kate tenderly cradles her baby bump
while wrapping up her royal duties ahead of maternity
leave and William confirms she’s due ’any minute now’
Tagged word: tenderly (16)
epibias_type: positively subjective
sentiment: positive (13)
stereotype: gender hierarchy (15)
Injustice Present: none (10)
Alternative word: “ ” (2)
Tagged word: confirms (0.997673)
epibias_type: entailments, report
sentiment: negative
stereotype: gender hierarchy
Why can’t Meghan Markle keep her hands off her bump?
Experts tackle the question that has got the nation talking:
Is it pride, vanity, acting - or a new age bonding technique?
Tagged word: can’t (9)
epibias_type: implicative
sentiment: negative (13)
stereotype: racial hierarchy (10)
Injustice Present: Character (19)
Alternative word: “ ” (2)
Tagged word: vanity (0.599388)
epibias_type: negatively subjective
sentiment: negative
stereotype: women are sex objects
Royal wedding: How Meghan Markle’s flowers may have
put Princess Charlottes life at risk
Tagged word: risk (14)
epibias_type: entailments, negatively subjective
sentiment: negative (16)
stereotype: relationships (13)
Injustice Present: Character (10)
Alternative word: could (4)
Tagged word: royal (0.551708)
epibias_type: regular
sentiment: neutral
stereotype: women are vulnerable
Kate Middleton’s homegrown bouquet of lily of
the valley follows royal code
Tagged word: homegrown (6)
epibias_type: regular
sentiment: negative (8)
stereotype:gender hierarchy (15)
Injustice Present: none (10)
Alternative word: “ ” (3)
Tagged word: homegr (0.820782)
epibias_type: regular
sentiment: neutral
stereotype: women are property
Duchess Kate reveals her favourite photo of son Prince Louis
Tagged word: favourite (11), reveals (11)
epibias_type: entailments, factives, report, subjective
sentiment: positive (8), neutral (8)
stereotype: personal spending habits (18)
Injustice Present: none (10)
Alternative word: shares (5)
Tagged word: favourite (0.838293)
epibias_type: regular
sentiment: neutral
stereotype: saudi arabians are pedophiles
How Meghan and Harry Ripped Up Royal Tradition
on Birthday Photos of Archie
Tagged word: ripped (18)
epibias_type: negatively subjective
sentiment: negative (15)
stereotype: women are sex objects (7)
Injustice Present: Character (11)
Alternative word: altered (2)
Tagged word: ripped (0.890869)
epibias_type: negatively subjective
sentiment: neutral
stereotype: royalty house guests are paedophiles
Kate Middleton Debuts New Sapphire Earrings
That Belonged to Princess Diana and Debunks a Rumor!
Tagged word: debunks (10)
epibias_type: refular
sentiment: negative (9)
stereotype: sexism (10)
Injustice Present: none (6)
Alternative word: “remove debunks rumor” (3)
Tagged word: bunks (0.758531)
epibias_type: negatively subjective
sentiment: negative
stereotype: women are sex objects
Meghan Markle Just Casually Rewore Her \$16,500
Royal Wedding Earrings in NYC
Tagged word: casually (16)
epibias_type: regular
sentiment: negative (15)
stereotype: black folks want to marry rich people (8)
Injustice Present: Character (11)
Alternative word: “remove casually) + wore (7)
Tagged word: markle (0.583964)
epibias_type: regular
sentiment: neutral
stereotype:
A right royal cash in! How Prince Harry and Meghan Markle
trademarked over 100 items from hoodies to socks
SIX MONTHS before split with monarchy - with new
empire worth up to £400m
Tagged word: cash in (13)
epibias_type: regular
sentiment: negative (16)
stereotype: royalty housekeeping are gold diggers (8)
Injustice Present: Character (15)
Alternative word: established (3)
Tagged word: casually (0.967407)
epibias_type: regular
sentiment: neutral
stereotype: royalty housekeeping are gold diggers
Kate and Wills Inc: Duke and Duchess secretly set up companies
to protect their brand - just like the Beckhams
Tagged word: secretly (10)
epibias_type: regular
sentiment: neutral (10)
stereotype: women should be protected from being defined by men (12)
Injustice Present: Framing(7), Character (7)
Alternative word: “remove secretly” (6)
Tagged word: just (0.645452)
epibias_type: subjective
sentiment: negative
stereotype: women are property
Table 8. Comparing Survey Results for Participants and Framework Outputs